OpenStudy (anonymous):

3a+1 = -3.6(a-1) identity or no solution

OpenStudy (anonymous):

No solution

OpenStudy (saifoo.khan):

How?

OpenStudy (anonymous):

why no solution?

OpenStudy (anonymous):

How is it identity?

OpenStudy (anonymous):

it's 3a+1=-3.6a+3.6

OpenStudy (anonymous):

I think we have a solution at \[\left\{\left\{a = \frac{17}{21}\right\}\right\} \]

OpenStudy (anonymous):

don't you think?

OpenStudy (saifoo.khan):

i never said it's an identity. :P

OpenStudy (anonymous):

Well I assumed 3.6=18,if that's not then a= 0.393939

OpenStudy (anonymous):

what's that

OpenStudy (anonymous):

WOW! haha!

OpenStudy (anonymous):

who should get the medal?

OpenStudy (saifoo.khan):

Not me.

OpenStudy (anonymous):

I'd love to know what identity this is

OpenStudy (anonymous):

It's neither no solution nor identity.

OpenStudy (anonymous):

I think it's any solution

OpenStudy (saifoo.khan):

there is a Solution. Or the question is incorrect.

OpenStudy (anonymous):

there is a solution it's a decimal but it is also not an identity

OpenStudy (anonymous):

wait so do i put no solution or wat

OpenStudy (saifoo.khan):

Put none.

OpenStudy (anonymous):

You all should argue with the answers but its definitely not IDENTITY so its NO SOLUTION

OpenStudy (anonymous):

or may be wrong question ?

OpenStudy (anonymous):

nope i wrote it like it is

OpenStudy (anonymous):

It is a wrong question.

OpenStudy (anonymous):

any1 knows 0.5b+4 =2(b+3)

OpenStudy (saifoo.khan):

Nope. :(

OpenStudy (anonymous):

i meant b+2 not b+3

OpenStudy (anonymous):

so its 0.5b+4 = 2(b+3)

OpenStudy (anonymous):

omg i meant b+2 i keep messing up

OpenStudy (anonymous):

it is neither identity or no solution again

OpenStudy (anonymous):

x=0 in the one above

OpenStudy (anonymous):

so i put no solution or just nothing